OpenStudy (anonymous):

PLEASE HELP !!! Which two numbers have a percent of increase of approximately 21% ?

pooja195 (pooja195):

Ok you need the difference between the two numbers, then you divide that by the number you started with. So in your case it would be Hope that helps.

OpenStudy (anonymous):

@pooja195 ,Thank you sooo muchh!:p

OpenStudy (anonymous):

But what are the 2 numbers,how do i get them???

OpenStudy (whpalmer4):

There are infinitely many pairs of numbers which have a percentage of increase of approximately 21%. There must be more to your problem. Or maybe you are just supposed to provide such a pair.

OpenStudy (anonymous):

Well, it just asks the questions and the choices to it.. a. 12 to 19 b. 101 to150 c. 10 to 21 d. 126 to152

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Okay, which of those pairs has one number that is 21% greater than the other?

OpenStudy (whpalmer4):

If you have two numbers, a and b, the percentage a > b is given by \[\frac{(a-b)}{b}*100\%\] So, if a = 3, b = 2, the percentage 3 is greater than 2 is \[\frac{3-2}{2}*100\% = 50\%\]

OpenStudy (whpalmer4):

If you want to find a number a that is x% larger than a number you have (b), \[a = (1+\frac{x}{100\%})*b\] If the number you have is 1000, and you want to find a number 21% greater: \[a = (1+\frac{21\%}{100\%})*1000=1.21*1000 = 1210\]

OpenStudy (whpalmer4):

So, look at the numbers in your answers: does 1.21* the first number give you the second number for any of them? Be careful, there are some trick answers in there.

OpenStudy (anonymous):

@whpalmer4 , So would it be D. , i did the (a-b/bx100%) and i got -26%

OpenStudy (whpalmer4):

Yes, it would be letter D, but I'm not sure I understand how you got -26%

OpenStudy (anonymous):

i got a. -7%, b.-49%, c. -5%, d.-25%

OpenStudy (anonymous):